# SplatSSC: Decoupled Depth-Guided Gaussian Splatting for Semantic Scene Completion

### v2

@@ Abstract
Monocular 3D Semantic Scene Completion (SSC) is a challenging yet promising task that aims to infer dense geometric and semantic descriptions of a scene from a single image. While recent object-centric paradigms significantly improve efficiency by leveraging flexible 3D Gaussian primitives, they still rely heavily on a large number of randomly initialized primitives, which inevitably leads to 1) inefficient primitive initialization and 2) outlier primitives that introduce erroneous artifacts. In this paper, we propose SplatSSC, a novel framework that resolves these limitations with a depth-guided initialization strategy and a principled Gaussian aggregator. Instead of random initialization, SplatSSC utilizes a dedicated depth branch composed of a Group-wise Multi-scale Fusion (GMF) module, which integrates multi-scale image and depth features to generate a sparse yet representative set of initial Gaussian primitives. To mitigate noise from outlier primitives, we develop the Decoupled Gaussian Aggregator (DGA), which enhances robustness by decomposing geometric and semantic predictions during the Gaussian-to-voxel splatting process. Complemented with a specialized Probability Scale Loss, our method achieves state-of-the-art performance on the Occ-ScanNet dataset, outperforming prior approaches by over $6.3\%$ in IoU and $4.1\%$ in mIoU, while reducing both latency and memory cost by more than $9.3\%$.
@@ Introduction
3D scene understanding has garnered significant attention with the rapid evolution of embodied agents (deng2025mcnslammultiagentcollaborativeneural; deng2025mne; deng2024plgslam) and autonomous driving (li2025flagdroneacing; yi2025embodiedgame; fusic2024improvedrrt; zheng2025robustefficient; zammit2023realtimeuav). As a key technology in this domain, 3D occupancy prediction (tong2023scene; huang2023tri; wei2023surroundocc; tian2023occ3d; wang2024panoocc) and 3D Semantic Scene Completion (SSC) (cao2024monoscene; miao2023occdepth; zhang2023occformer; li2023voxformer; mei2024camera) have made remarkable progress. Early and conventional approaches for these tasks predominantly rely on grid-based representations. However, processing dense 3D volumes incurs prohibitive computational and memory costs. To mitigate this limitation, various efficiency-driven strategies have been explored, such as accelerating processing with Bird’s-Eye-View (BEV) projections (yu2023flashocc; hou2024fastocc), or leveraging the natural sparsity of scenes with sparse voxels (tang2024sparseocc; li2023voxformer) and points (shi2024occupancy; wang2024opus).
@@
Figure 3: Illustration of the proposed GCA layer. The weight matrix $W_{a}$ is shared across different groups and scales, thus reducing memory consumption and computational cost to obtain the attention score.
@@
where $W_{q}$, $W_{k}$, and $W_{v}$ are linear projection matrices for Query, Key, and Value, respectively. $l\in\{1,...,L\}$ denotes the scale index. Inspired by the efficient design of Deformable Attention (zhu2020deformable), we adopt a lightweight linear projection mechanism in place of the standard dot-product attention. To elaborate, the attention scores are computed by feeding the element-wise sum of queries and keys into a shared projection $W_{a}\in\mathbb{R}^{D_{g}\times 1}$
@@
The design of GCA is computationally lean. Standard cross-attention has a complexity of $\mathcal{O}(LN^{2}D)$, where $N$ is the sequence length. In contrast, by employing a group-wise mechanism and replacing the quadratic-cost dot-product with a linear-cost MLP, GCA significantly reduces the complexity. The dominant cost of our module becomes $\mathcal{O}(ND^{2}(L+2)/G)$, which is substantially more efficient, especially for long feature sequences.
@@ Acknowledgments
We provide additional qualitative visualizations of our depth branch in Figure 8. For each scene, we show the ground-truth depth at its naively downsampled $30\times 40$ version, followed by the output of our depth branch and FT-DAv2 at the same $30\times 40$ resolution, which can be used to initialize the initial Gaussian primitives in location.
Compared to FT-DAv2, our depth branch produces depth maps that are visibly closer to the ground truth. It better maintains a clearer separation between foreground objects and background surfaces.

### v3

@@ Abstract
Monocular 3D Semantic Scene Completion (SSC) is a challenging yet promising task that aims to infer dense geometric and semantic descriptions of a scene from a single image. While recent object-centric paradigms significantly improve efficiency by leveraging flexible 3D Gaussian primitives, they still rely heavily on a large number of randomly initialized primitives, which inevitably leads to 1) inefficient primitive initialization and 2) outlier primitives that introduce erroneous artifacts. In this paper, we propose SplatSSC, a novel framework that resolves these limitations with a depth-guided initialization strategy and a principled Gaussian aggregator. Instead of random initialization, SplatSSC utilizes a dedicated depth branch composed of a Group-wise Multi-scale Fusion (GMF) module, which integrates multi-scale image and depth features to generate a sparse yet representative set of initial Gaussian primitives. To mitigate noise from outlier primitives, we develop the Decoupled Gaussian Aggregator (DGA), which enhances robustness by decomposing geometric and semantic predictions during the Gaussian-to-voxel splatting process. Complemented with a specialized Probability Scale Loss, our method achieves state-of-the-art performance on the Occ-ScanNet dataset, outperforming prior approaches by over 6.3% in IoU and 4.1% in mIoU, while reducing both latency and memory cost by more than 9.3%.
@@ Introduction
3D scene understanding has garnered significant attention with the rapid evolution of embodied agents and autonomous driving (li2025flagdroneacing; yi2025embodiedgame; fusic2024improvedrrt; zheng2025robustefficient; zammit2023realtimeuav). As a key technology in this domain, 3D occupancy prediction (tong2023scene; huang2023tri; wei2023surroundocc; tian2023occ3d; wang2024panoocc) and 3D Semantic Scene Completion (SSC) (cao2024monoscene; miao2023occdepth; zhang2023occformer; li2023voxformer; mei2024camera) have made remarkable progress. Early and conventional approaches for these tasks predominantly rely on grid-based representations. However, processing dense 3D volumes incurs prohibitive computational and memory costs. To mitigate this limitation, various efficiency-driven strategies have been explored, such as accelerating processing with Bird’s-Eye-View (BEV) projections (yu2023flashocc; hou2024fastocc), or leveraging the natural sparsity of scenes with sparse voxels (tang2024sparseocc; li2023voxformer) and points (shi2024occupancy; wang2024opus).
@@
Figure 3: Illustration of the proposed GCA layer. The weight matrix $W_{a}$ is shared across different groups and scales, thus reducing memory consumption and computational cost.
@@
where $W_{q}$, $W_{k}$, and $W_{v}$ are linear projection matrices for Query, Key, and Value, respectively. $l\in\{1,...,L\}$ denotes the scale index. Inspired by the efficient design of Deformable Attention (zhu2020deformable), we adopt a lightweight linear projection mechanism in place of the standard dot-product attention. To elaborate, the attention scores are computed by feeding the element-wise sum of queries and keys into a shared projection $W_{a}\in\mathbb{R}^{D_{g}\times 1}$:
@@
The design of GCA is computationally lean. Standard cross-attention has a complexity of $\mathcal{O}(LN^{2}D)$, where $N$ is the sequence length. In contrast, by employing a group-wise mechanism and replacing the quadratic-cost dot-product with a linear-cost MLP, GCA significantly reduces the complexity. The dominant cost of our module becomes $\mathcal{O}(ND^{2}(L+2)/G)$, which is substantially more efficient, especially for long sequences.
@@ Acknowledgments
Figure 8 provides additional qualitative comparisons of our depth branch. We visualize $30\times 40$ depth maps from our method and FT-DAv2 against downsampled ground truth, which serve as spatial priors for Gaussian primitive initialization. Compared to FT-DAv2, our branch produces depth maps with higher fidelity and sharper foreground-background separation, aligning more closely with the ground truth.